OpenStudy (anonymous):

Solve for x. 2x2 - 4x = 0 a. 0, -4 b. 0, -2 c. 0, 2 d. 2, 4 help me please <3

OpenStudy (amoodarya):

\[2x^2-4x=0\\2(x^2-2x)=0\\2x(x-2)=0\\\]

OpenStudy (anonymous):

hu? can you explain this please?

OpenStudy (amoodarya):

factoring simplifying

OpenStudy (amoodarya):

when A.B=0 A=0 or B=0 ok ?

OpenStudy (anonymous):

damn i dont know >.< please keep helping me.

OpenStudy (amoodarya):

if x-3=0 what is x ? second question : when x (x-3) =0 what are value of x ?

OpenStudy (anonymous):

okay i got it its c :)

OpenStudy (amoodarya):

ok

OpenStudy (amoodarya):

you are right

OpenStudy (anonymous):

thanks :D